If c does not equal 0 and the slope of the line passing through (-c, 2c) and (4c, a) is 1, what is the expression for a in terms of c? Hint: Set up the slope formula, substitute, and solve the resulting proportion (cross multiplication).


Answer by ilana(307)

the slope of a line going through points (x1,y1) and (x2,y2) is (y2-y1)/(x2-x1), so in this case just substitute in the coordinates to make an equation with variables. (a-2c)/(4c--c)=1
(a-2c)/(4c+c)=1
(a-2c)/5c=1
Cross multiply (or multiply both sides of the equation by 5c) to get a-2c=5c.
You want to solve for a, so add 2c to both sides: a=7c. That is your solution.